OpenStudy (anonymous):

The formula Q = MCT where Q = heat flow, M = mass, C = specific heat, and T = change of temperature is used to calculate heat flow. Solve this formula for T.

Vocaloid (vocaloid):

"solve for T" means isolate T divide both sides by M and C

Vocaloid (vocaloid):

try and see what happens

OpenStudy (anonymous):

Solve 2(x + 1) = 2x + 5.

OpenStudy (anonymous):

Solve −2(2x + 5) − 3 = −3(x − 1).